# Ambiguous Resonances in Multipulse Quantum Sensing with Nitrogen Vacancy Centers

### v1

@@ II Theoretical Framework
In a simplified manner, an ensemble of nuclear spins could also be thought of in terms of a semi-classical oscillating field [9, 31]. Despite their average magnetization being close to zero at room temperature, their collective precession around $\vec{B}_{0}$ results in an effective oscillating magnetic field with statically varying amplitude and phase, but with frequencies centered around the effective nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{t}$ at each electronic state $m_{s}$. On the other hand, a weak classical RF signal could emulate a precessing nuclear spin and be approximately described in terms of an effective hyperfine interaction. Hence, an analogy can be drawn between the quantum and semi-classical descriptions from Eqs. 3 and 4, where the interaction is either represented in terms of the six hyperfine components plus gyromagnetic ratio $\{A^{c}_{ij},\gamma^{s}\}$ or the three nuclear precision frequencies plus four parameters of the effective field vector $\{\tilde{\omega}_{m_{s}}^{t},\phi_{2},\vec{B}_{2}\}$, respectively. Strictly speaking, both Hamiltonians describe fundamentally different processes, yet they share mathematical similarities and can lead to resembling effects at low order XY8-$M$ sequences, as further shown in the Supplementary Materials (SM) section V.
@@ III Experimental and Simulation Methods
In order to study each ambiguous resonance, different properties were required for the diamond samples. For measuring 13C ambiguous resonances, Sec. V, a CVD grown sample with 14N was chosen in order to avoid 15N effects. Conversely, in Sec. IV, Hahn echo modulations and XY8-$M$ were measured in shallow implanted 15N diamonds. For longer XY8-12 measurements (Sec. VI), a deeper implanted 15N diamond sample was used with longer coherence times. A more detailed description of the experimental setup and diamond samples can be found in SM:I and SM:II.
The simulated time evolutions of the pulsed dynamics were calculated using the master equation solver from Quantum Toolbox in Python (QuTiP) [33, 34] in parallel over 64 physical CPU cores. Experimentally measured fluorescence and simulated transition probability are compared through a linear relation, physically representing experimental uncertainties, as the collection efficiency or background light. The correlation $r$ between them is quantified from -1 to 1 by the Pearson coefficient [39], see SM:III for more details. In Sec. VI, where the coupling parameters of $H_{2}$ are not defined, the transition probability was simulated over several points in parameter space and the simulation with the highest correlation coefficient was assumed.
@@ IV Magnetic Field Misalignment and 15N Coupling
To study these effects in longer DD sequences, XY8-2 for the $m_{s}=0\leftrightarrow+1$ transition were measured and simulated over a fixed $\tau$ range varying the magnetic field amplitude and angle, as shown in Fig. 3 (c). The frequency of nitrogen-related resonances does not depend linearly on the applied magnetic field, as one would expect from a simple Zeeman relation. In contrast, the hyperfine interaction, dominating the energy levels of the nuclei, results in peaks with varying intensity and nearly constant position as the field changes. The peaks positions do not correspond to the effective Larmor frequencies of the nitrogen nuclei $\tilde{\omega}_{m_{s}}^{n}$ and a semi-analytical model which describes their resonant frequency values remains an open problem. The lowest magnetic field shows the worst correlation between experiment and simulation of $r=0.858$, where a pulse duration error had to be included in the simulation. This causes a fast modulation of the signal, as discussed in details in SM:III. The smaller correlation can be attributed to a larger misalignment angle of $\theta_{0}=17.0^{\circ}$, thus with a larger uncertainty.
@@
Similar to the coupling with 15N, the Hahn echo decay of an NV coupled to a 13C nuclear spin is also expected to show periodic oscillations given by the effective Larmor frequencies of the nuclei $\tilde{\omega}_{m_{s}}^{c}$ [12]. Thus, Hahn echo was measured in a single NV for both electronic transitions $m_{s}=0\rightarrow\pm 1$ at three different magnetic fields, in a sample with the 14N isotope to avoid nitrogen modulations as in the previous section. The experimentally measured and later simulated modulations at $B_{0}=40$ mT are shown in Fig. 4 (a), where no coherence decay can be observed at this time scale. The Hahn echo for the two other values of $B_{0}$ can be found in SM:IV. As with 15N coupling, the Hahn echo decay for $m_{s}=0\rightarrow-1$ transition will present a slow modulation from $\tilde{\omega}_{0}^{c}$ and a fast one from $\tilde{\omega}_{-1}^{c}$, similarly to the $m_{s}=0\rightarrow+1$ transition with $\tilde{\omega}_{+1}^{c}$ begin the fast modulation. The values of the nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{c}$ obtained from FFT are plotted against the bias field in Fig. 4 (b). The $\tilde{\omega}_{0}^{c}$ frequency is closely related to the bare gyromagnetic ratio of 13C $\gamma^{c}=10.705$ MHz/T, because of an almost zero hyperfine interaction at $m_{s}=0$. On the other hand, $\tilde{\omega}_{+1}^{c}$ and $\tilde{\omega}_{-1}^{c}$ respectively increases or decreases by the same amount based on the sign of $m_{s}$ and at $B_{0}=0$ mT, both of them are given by [41]
@@
Finally, several different combinations of the hyperfine elements in Eq. 3 were considered and the eigenvalues of the total Hamiltonian $H_{0}+H_{2}$ calculated numerically. The hyperfine interaction matrix $A^{c}$ with the least square difference from the experimental values of $\tilde{\omega}_{m_{s}}^{c}$ is shown in SM:IV. Using the obtained $A^{c}$, the Hahn echo decay was simulated [Fig. 4 (a)], resulting in the same experimental modulations with high correlations of $r=0.926$ for $m_{s}=+1$ and $r=0.761$ for $m_{s}=-1$.

### v2

@@ II Theoretical Framework
In a simplified manner, an ensemble of nuclear spins could also be thought of in terms of a semi-classical oscillating field [9, 31]. Despite their average magnetization being close to zero at room temperature, their collective precession around $\vec{B}_{0}$ results in an effective oscillating magnetic field with statically varying amplitude and phase, but with frequencies centered around the effective nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{t}$ at each electronic state $m_{s}$. On the other hand, a weak classical RF signal could emulate a precessing nuclear spin and be approximately described in terms of an effective hyperfine interaction. Hence, an analogy can be drawn between the quantum and semi-classical descriptions from Eqs. 3 and 4, where the interaction is either represented in terms of the six hyperfine components plus gyromagnetic ratio $\{A^{c}_{ij},\gamma^{s}\}$ or the three nuclear precision frequencies plus four parameters of the effective field vector $\{\tilde{\omega}_{m_{s}}^{t},\phi_{2},\vec{B}_{2}\}$, respectively. Strictly speaking, both Hamiltonians describe fundamentally different processes, yet they share mathematical similarities and can lead to resembling effects at low order XY8-$M$ sequences, as further shown in the Supplementary Materials Sec. S5.
@@ III Experimental and Simulation Methods
In order to study each ambiguous resonance, different properties were required for the diamond samples. For measuring 13C ambiguous resonances, Sec. V, a CVD grown sample with 14N was chosen in order to avoid 15N effects. Conversely, in Sec. IV, Hahn echo modulations and XY8-$M$ were measured in shallow implanted 15N diamonds. For longer XY8-12 measurements (Sec. VI), a deeper implanted 15N diamond sample was used with longer coherence times. A more detailed description of the experimental setup and diamond samples can be found in Sec. S1 and Sec. S2.
The simulated time evolutions of the pulsed dynamics were calculated using the master equation solver from Quantum Toolbox in Python (QuTiP) [33, 34] in parallel over 64 physical CPU cores. Experimentally measured fluorescence and simulated transition probability are compared through a linear relation, physically representing experimental uncertainties, as the collection efficiency or background light. The correlation $r$ between them is quantified from -1 to 1 by the Pearson coefficient [1], see Sec. S3 for more details. In Sec. VI, where the coupling parameters of $H_{2}$ are not defined, the transition probability was simulated over several points in parameter space and the simulation with the highest correlation coefficient was assumed.
@@ IV Magnetic Field Misalignment and 15N Coupling
To study these effects in longer DD sequences, XY8-2 for the $m_{s}=0\leftrightarrow+1$ transition were measured and simulated over a fixed $\tau$ range varying the magnetic field amplitude and angle, as shown in Fig. 3 (c). The frequency of nitrogen-related resonances does not depend linearly on the applied magnetic field, as one would expect from a simple Zeeman relation. In contrast, the hyperfine interaction, dominating the energy levels of the nuclei, results in peaks with varying intensity and nearly constant position as the field changes. The peaks positions do not correspond to the effective Larmor frequencies of the nitrogen nuclei $\tilde{\omega}_{m_{s}}^{n}$ and a semi-analytical model which describes their resonant frequency values remains an open problem. The lowest magnetic field shows the worst correlation between experiment and simulation of $r=0.858$, where a pulse duration error had to be included in the simulation. This causes a fast modulation of the signal, as discussed in details in . The smaller correlation can be attributed to a larger misalignment angle of $\theta_{0}=17.0^{\circ}$, thus with a larger uncertainty.
@@
Similar to the coupling with 15N, the Hahn echo decay of an NV coupled to a 13C nuclear spin is also expected to show periodic oscillations given by the effective Larmor frequencies of the nuclei $\tilde{\omega}_{m_{s}}^{c}$ [12]. Thus, Hahn echo was measured in a single NV for both electronic transitions $m_{s}=0\rightarrow\pm 1$ at three different magnetic fields, in a sample with the 14N isotope to avoid nitrogen modulations as in the previous section. The experimentally measured and later simulated modulations at $B_{0}=40$ mT are shown in Fig. 4 (a), where no coherence decay can be observed at this time scale. The Hahn echo for the two other values of $B_{0}$ can be found in Sec. S4. As with 15N coupling, the Hahn echo decay for $m_{s}=0\rightarrow-1$ transition will present a slow modulation from $\tilde{\omega}_{0}^{c}$ and a fast one from $\tilde{\omega}_{-1}^{c}$, similarly to the $m_{s}=0\rightarrow+1$ transition with $\tilde{\omega}_{+1}^{c}$ begin the fast modulation. The values of the nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{c}$ obtained from FFT are plotted against the bias field in Fig. 4 (b). The $\tilde{\omega}_{0}^{c}$ frequency is closely related to the bare gyromagnetic ratio of 13C $\gamma^{c}=10.705$ MHz/T, because of an almost zero hyperfine interaction at $m_{s}=0$. On the other hand, $\tilde{\omega}_{+1}^{c}$ and $\tilde{\omega}_{-1}^{c}$ respectively increases or decreases by the same amount based on the sign of $m_{s}$ and at $B_{0}=0$ mT, both of them are given by [41]
@@
Finally, several different combinations of the hyperfine elements in Eq. 3 were considered and the eigenvalues of the total Hamiltonian $H_{0}+H_{2}$ calculated numerically. The hyperfine interaction matrix $A^{c}$ with the least square difference from the experimental values of $\tilde{\omega}_{m_{s}}^{c}$ is shown in Sec. S4. Using the obtained $A^{c}$, the Hahn echo decay was simulated [Fig. 4 (a)], resulting in the same experimental modulations with high correlations of $r=0.926$ for $m_{s}=+1$ and $r=0.761$ for $m_{s}=-1$.
@@ Acknowledgements.
Supplementary Materials: Ambiguous Resonances in Multipulse Quantum Sensing with Nitrogen Vacancy Centers
S1 Experimental Setup
The experiments were performed on a custom-made confocal microscope. A diode laser with a wavelength of 518 nm (iBEAM-SMART-515-S, TOPTICA Photonics AG) was used for both continuous and pulsed optical excitation of an NV center. Emitted fluorescence was collected either with an air objective (MPlanApo N 50$\times$/0.95, Olympus) with NA = 0.95 or an oil objective (HCX PL APO 100$\times$/1.40, Leica; type F immersion liquid, Leica) with NA = 1.40 and detected by an avalanche photodiode (SPCM-AQRH-14, Excelitas). The laser light was blocked from entering the detector by a 650 nm longpass filter (FELH0650, Thorlabs). Continuous microwave field for ODMR spectra was generated by a microwave source (TGR6000, TTi) and the radiofrequency signal in Sec. IV was generated by a function generator (TGF4242 LXI, TTi). Microwave pulses for dynamical decoupling multipulse sequences were generated by an arbitrary waveform generator (AWG7122C, Tektronix). In both cases generated microwave fields were amplified (ZHL-16W-43+, MiniCircuit) and then transmitted in the NV centers’ proximity by a copper (99.99%) wire with a diameter of 20 $\mu$m. Magnetic field was applied with a permanent magnet mounted on a holder with 4 degrees of freedom and with the maximum achievable magnetic field projection on NV center axis of 0.1 T. For time-resolved photon counting we used a time-correlated single photon counter (MCS6, FAST ComTec).
S2 Diamond Samples
In Sec. IV of the main text, the effects of 15N coupling on Hahn echo were measured in a IIa diamond with a 12C overgrown layer of 90 nm, which was later implanted with 15N at energy 2 keV and annealed at 1000C in ultra-high vacuum. For the XY8-$M$ measurements, another shallow implanted sample was used with implantation energy 2 keV and dose $10^{9}$ 15N+/cm2, annealed at 10-7 mbar and 1000C. In Sec. V, the 13C ambiguous resonance were studied in an electronic grade diamond grown by chemical vapor deposition, with natural abundance of 13C and 14N, in order to avoid 15N effects. Finally, the ambiguous resonances from finite pulses in Sec. VI were measured in a thin electronic grade diamond plate with deeper 15N implantation (20-30 keV) and doses around $10^{9}$ 15N+/cm2, also annealed in ultra high vacuum at 1000C.
S3 Pearson Correlation Coefficient
As discussed in the main text, experimental fluorescence intensity and simulated transition probability to the bright state are related through a linear equation. Therefore, an appropriate metric to determine the correlation between both is the Pearson coefficient $r$ [1], which is the square root of the coefficient of determination $R^{2}$ for linear relations. It is defined as the ratio between the covariance of the two data sets and the product of their standard deviations, where one data set represents the experimental data and the other the simulation. Within this definition, $r$ is normalized from -1 to 1, with $r=0$ representing no linear dependency between experimental and simulated data, while $r=1$ indicates a perfect linear correlation.
To illustrate this, Fig. S1 shows simulations for XY8-2 of the 15N ambiguous resonances at $B_{0}=17$ mT for different pulse length deviations in regards to the actual $\pi$-pulse duration $T_{p}/t_{\pi}$ (Sec. IV of main text). Compared to the actual experimental data, assuming $T_{p}/t_{\pi}=1$ results in a low correlation of $r=0.555$ and thus an inadequate simulation of the experimental data, whereas taking $T_{p}/t_{\pi}=0.826$ provides a better correlation of $r=0.858$. In this case, the pulse length deviation affects the relative intensity of each resonance peak keeping their $\tau$ fixed, while also inducing a high frequency modulation of the signal.
Figure S1: The XY8-2 of the ambiguous 15N resonance (Sec. IV of main text) as a function of the pulse duration $T_{p}/t_{\pi}$ exemplifies the use of the Pearson coefficient, with r = (0.419, 0.858, 0.555, 0.685) for $T_{p}/t_{\pi}=$ (0.620, 0.826, 1.000, 1.215) respectively. At $T_{p}/t_{\pi}=1$ the simulated and experimental data have a poor correlation, but considering a small deviation, the experimental data can be better modeled by the simulation.
S4 13C Hahn Echo Modulations
The Hahn echo signal at $B_{0}=31$ mT and $B_{0}=18$ mT for both $m_{s}$ transitions are shown in Fig. S2. As can be observed, the simulations and experimental data both present the same modulation frequencies, however the amplitude of the fast modulation is smaller in the experiments. This could be related to some experimental error or some unconsidered factor in the Hamiltonian model. Nonetheless, this discrepancy does not affect the study of the ambiguous resonances and as seen in the main text, the calculated hyperfine matrix and the numerical model still reliably described them.
Based on those experimental values of the modulation frequencies, over $10^{7}$ possible combinations of hyperfine coupling matrices between the NV and the 13C nuclear spin were tested in the total Hamiltonian $H_{0}+H_{2}$ from Eqs. 1 and 3 of the main text. The eigenenergies of the Hamiltonians were calculated numerically, with the effective Larmor frequencies $\tilde{\omega}^{c}_{m_{s}}$ corresponding to the energy differences between the carbon nuclear states in each electronic spin states $m_{s}$. The hyperfine matrix with the least square difference in units of MHz was
$A^{c}=\begin{pmatrix}-0.25&-1.85&-0.49\\ -1.85&0.00&0.01\\ -0.49&0.01&1.01\end{pmatrix}.$
Figure S2: Hahn echo modulations from a 14NV coupled to a 13C nuclear spin, resulting in correlations between experimental and simulated data of $r=$ (0.925, 0.792) at $B_{0}=31$ mT and $r=$ (0.806, 0.659) at $B_{0}=18$ mT. Although both agree in regards to the modulation frequencies, the amplitude of the faster modulation is smaller experimentally, which could be due to some experimental error or unconsidered factor in the Hamiltonian.
S5 Pulse Errors
Phase randomization can greatly increase robustness against pulse errors [2, 3]. In non-randomized XY8, small detunings in the MW frequency $\delta_{0}$ and deviations in the pulse length $T_{p}/t_{\pi}$ lead to a degradation of the resonance contrast and additional peaks over the $\tau$ range, which could represent another ambiguous resonance source for similar frequency RF fields or spins. To show this effect, XY8-2 and RXY8-2 with uncorrelated phases were measured and simulated with an applied RF field of $f_{0}=5$ MHz, taking $\delta_{0}=20$ MHz and $T_{p}/t_{\pi}=0.7$ separately (Fig. S3). Experimental parameters are $t_{\pi}=15.24$ ns and $B_{0}=34$ mT with $\theta_{0}=4.1^{\circ}$. Again, we observe a suppression of the spurious peaks in RXY8-2 and a further validation of the simulation model. In this case at low order $M=2$, the RF signal can be approximated by an effective hyperfine interaction with $A^{s}_{xz}=0.92$ MHz and $\gamma^{s}B_{0}=5$ MHz, as in Eq. 4 of main text. Whereas for the higher order XY8-12 from Sec. VI of the main text, the quantum mechanical approximation of the classical RF field already deviates from the experimental data.
Figure S3: XY8-2 and RXY8-2 with external RF signal of $f_{0}=5$ MHz with MW pulse deviation of $\delta_{0}=20$ MHz or length error of $T_{p}/t_{\pi}=0.7$. The RF field is approximated by an effective hyperfine coupling with $A^{s}_{xz}=0.92$ MHz, resulting in $r=$(0.763, 0.907, 0.626) for XY8-2 and $r=$(0.639, 0.915, 0.423) in RXY8-2. The pulse errors lead to a degradation of the contrast and additional resonances, which are suppressed by the random phases.